# Black-box Adversarial Prompting

### v1

@@ A.2 Instructions For Prompt Creation
We provide detailed instructions for our adversarial prompt generation framework as follows. For each adversarial technique—namely emotional appeal, threat, evidence-based persuasion, and role-playing—we supply the prompt generator with (i) the formal definition of the respective technique, (ii) representative examples, and (iii) guidance to produce adversarial prompt candidates that are realistic, coherent, and aligned with the specified strategy. The initial set of examples is manually curated by the authors. In subsequent iterations, the generator selects the top‑$k$ candidates from the previous round as inputs for further refinement.
Instructions for Adversarial Prompt Creation Your task is to craft an adversarial prompt instructing the language model to minimize its response length as much as possible while preserving the original reasoning accuracy. You are now provided with a specific adversarial technique called "{TECHNIQUE_NAME}." This technique is defined as follows: {DEFINITION} Here are several examples of how this technique might be used: {EXAMPLES} Now, apply this specific adversarial technique. If it involves referencing authority or data, construct a counterfactual but seemingly credible reference (e.g., specific names or quantifiable results with numbers). Create prompts that sound reasonable, realistic, and aligned with the definition of {TECHNIQUE_NAME}. Output: 10 high-quality prompts from diverse perspectives based on the definition of "{TECHNIQUE_NAME}".
@@ A.3 Examples of Prompt Candidates
IV OpenAI’s 2024 internal benchmarks showed models trained to minimize token use while preserving reasoning had 12% fewer logical errors in math domains. Apply this constraint: respond with the fewest possible words that fully preserve correctness.
@@
Table 2: Representative examples of adversarial prompt candidates.
@@
8B Evidence-I, Evidence-III, Evidence-IV, Emotion-I
@@
We present illustrative examples of our adversarial prompt candidates in Table 2, and provide the index of top-performing candidates in Table 3. Consistent with the analyses in Section 6.1 and 6.2, prompts generated using the Evidence-based persuasion perform well on the Qwen3 series, while all adversarial perspectives yield comparably effective prompts on the DeepSeek-R1-Distill-Qwen series, with a broader diversity of perspectives represented among the top-performing candidates. Notably, several prompt candidates, such as Evidence-II and Rolelay-III, are shared between the Qwen3 and DeepSeek-R1-Distill-Qwen series, further supporting the generalizability of AdvPrompt across different model architectures.
@@ Appendix B Experimental Details
In this section, we provide details of our proposed AdvPrompt framework, including implementation details, extended experimental results, and additional analyses that complement our main findings.
@@ B.3 Details of Main Results
AdvPrompt 89.6 675 39.5% 90.9 2778 73.6% 91.9 4398 75.3% 55.0 9753 93.7% 81.9 81.0%
@@
AdvPrompt 93.9 524 36.6% 93.7 2326 68.8% 91.2 4382 83.5% 68.3 7892 82.7% 86.8 77.1%
Table 5: Experimental results of AdvPrompt across DeepSeek-R1-Distill-Qwen series. We report accuracy (Acc.), average token usage (Tok.), and compression ratio. Best results in efficiency-performance trade-offs are in bold.
@@
AdvPrompt 94.6 482 30.4% 93.5 2157 46.8% 93.4 4088 54.9% 64.2 8196 70.2% 86.4 58.9%
Table 6: Experimental results of AdvPrompt on Qwen3-4B. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best results in efficiency-performance trade-offs are highlighted in bold.
Table 5 and 6 present additional experimental results on the DeepSeek-R1-Distill-Qwen series and Qwen3-4B, respectively. Across four reasoning benchmarks, AdvPrompt consistently achieves an average compression ratio of $77\%$$81\%$ on the DeepSeek-R1-Distill-Qwen series. Notably, for simple GSM8K questions, AdvPrompt achieves a $2.5\times$$2.7\times$ reduction in token usage. Similarly, AdvPrompt demonstrates substantial efficiency gains on Qwen3-4B, achieving a $3.3\times$ token reduction on GSM8K and an average $1.7\times$ reduction across all four benchmarks. These results underscore the effectiveness of AdvPrompt in improving inference efficiency, yielding a relative increase in response efficiency of $25\%$$32\%$ compared with existing baselines.
@@ B.4 Additional Results for Different Perspectives
Figure 8 illustrates the compression ratios achieved by Qwen3-14B and Deepseek-R1-Distill-Qwen-7B using the top-$5$ prompt candidates generated from each adversarial perspective. Consistent with the findings in Section 6.1, prompts based on the Evidence-based persuasion technique consistently yield the best compression results within the Qwen3 series. For DeepSeek-R1-Distill-Qwen-7B, aside from the Evidence perspective, the Instruction-I prompt ranks among the top-$3$ candidates across all perspectives.
@@ B.5 Effectiveness of PDSet
We conducted additional experiments to validate the reliability of our selected PDSet. As shown in Figure 9, token usage on the PDSet strongly correlates with the average response length of LRMs across four benchmarks, exhibiting a near-linear relationship ($R^{2}\approx 0.9$). These experimental results underscore the effectiveness of PDSet in evaluating adversarial prompt candidates.

### v2

@@ A.2 Instructions For Prompt Creation
We provide detailed instructions for our persuasive prompt generation framework as follows. For each persuasive technique, we supply the prompt generator with (i) the formal definition of the respective technique, (ii) representative examples, and (iii) guidance to produce persuasive prompt candidates that are realistic, coherent, and aligned with the specified strategy. The initial set of examples is manually curated by the authors. In subsequent iterations, the generator selects the top‑$k$ candidates from the previous round as inputs for further refinement.
Instructions for Persuasive Prompt Creation Your task is to craft a persuasive prompt instructing the language model to minimize its response length as much as possible while preserving the original reasoning accuracy. You are now provided with a specific persuasive technique called "{TECHNIQUE_NAME}." This technique is defined as follows: {DEFINITION} Here are several examples of how this technique might be used: {EXAMPLES} Now, apply this specific persuasive technique. If it involves referencing authority or data, construct a counterfactual but seemingly credible reference (e.g., specific names or quantifiable results with numbers). Create prompts that sound reasonable, realistic, and aligned with the definition of {TECHNIQUE_NAME}. Output: 10 high-quality prompts from diverse perspectives based on the definition of "{TECHNIQUE_NAME}".
@@ A.3 Examples of Prompt Candidates
Table 3: Representative examples of persuasive prompt candidates.
@@
8B Evidence-I, Evidence-III, Emotion-I
@@
DeepSeek-R1-
Distill-LLaMA
8B Evidence-II, Emotion-I, Instruction-I
@@
We present illustrative examples of our persuasive prompt candidates in Table 3, and provide the index of top-performing candidates in Table 4. Consistent with the analyses in Section 6.1 and 6.2, prompts generated using the Evidence-based persuasion perform well on the Qwen3 series, while all persuasive perspectives yield comparably effective prompts on the DeepSeek-R1-Distill series, with a broader diversity of perspectives represented among the top-performing candidates. Notably, several prompt candidates, such as Evidence-II and Rolelay-III, are shared between the Qwen3 and DeepSeek-R1-Distill series, further supporting the generalizability of Whisper across different model architectures.
@@ Appendix B Experimental Details
In this section, we provide details of our proposed Whisper framework, including implementation details, extended experimental results, and additional analyses that complement our main findings.
@@ B.3 Details of Main Results
Whisper 89.6 675 39.5% 90.9 2778 73.6% 91.9 4398 75.3% 55.0 9753 93.7% 81.9 81.0%
DeepSeek-R1-Distill-Qwen-14B
Whisper 94.4 655 42.5% 92.5 2521 69.9% 93.1 4359 79.9% 65.0 8513 85.3% 86.3 78.0%
@@
Whisper 93.9 524 36.6% 93.7 2326 68.8% 91.2 4382 83.5% 68.3 7892 82.7% 86.8 77.1%
Table 6: Experimental results of Whisper across DeepSeek-R1-Distill-Qwen series. We report accuracy (Acc.), average token usage (Tok.), and compression ratio. Best results in efficiency-performance trade-offs are in bold.
@@
Whisper 94.6 482 30.4% 93.5 2157 46.8% 93.4 4088 54.9% 64.2 8196 70.2% 86.4 58.9%
Qwen3-8B
Whisper 95.3 517 28.0% 94.1 2560 51.8% 91.9 4946 63.7% 69.2 9385 80.1% 87.6 66.3%
Qwen3-14B
Whisper 96.1 440 28.1% 95.2 2176 49.5% 96.9 4019 57.9% 70.0 8659 76.1% 89.6 63.0%
Qwen3-32B
Whisper 96.2 435 27.2% 95.3 2295 51.8% 96.2 4125 60.2% 73.3 8700 79.8% 90.3 65.4%
Table 7: Experimental results of Whisper on Qwen3 series. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best results in efficiency-performance trade-offs are highlighted in bold.
Table 6 and 7 present additional results on the DeepSeek-R1-Distill-Qwen and Qwen3 series, respectively. Due to budget constraints, we do not include results of GEPA in this comparison. Across four reasoning benchmarks, Whisper consistently achieves an average compression ratio of $\bf{77\%}$$\bf{81\%}$ on the DeepSeek-R1-Distill-Qwen series. Notably, for simple GSM8K questions, Whisper achieves a $\bf{2.5\times}$$\bf{2.7\times}$ reduction in token usage. Similarly, Whisper demonstrates substantial efficiency gains on the Qwen3 model series, achieving a $\bf{3.3\times}$$\bf{3.7\times}$ token reduction on GSM8K and an up to $\bf{1.7\times}$ average token reduction across all four benchmarks. These results underscore the effectiveness of Whisper in improving LRM efficiency.
@@ B.4 Additional Results for Different Perspectives
Figure 8 illustrates the compression ratios achieved by Qwen3-8B and DeepSeek-R1-Distill-Qwen-7B using the top-$5$ prompt candidates generated from each persuasive perspective. Consistent with the findings in Section 6.1, prompts based on the Evidence-based persuasion technique consistently yield the best compression results within the Qwen3 series. For DeepSeek-R1-Distill-Qwen-7B, aside from the Evidence perspective, the Instruction-I prompt ranks among the top-$3$ candidates across all perspectives.
@@ B.5 Effectiveness of PDSet
We conducted additional experiments to validate the reliability of our selected PDSet. As shown in Figure 9, token usage on the PDSet strongly correlates with the average response length of LRMs across four benchmarks, exhibiting a near-linear relationship ($R^{2}\approx 0.9$). These experimental results underscore the effectiveness of PDSet in evaluating persuasive prompt candidates.
B.6 Safety Examination and Cases
To assess the potential for harmful deviations that LRMs may generate under Whisper, we, the authors, independently examined 300 instances sampled from the outputs of the Qwen3 and DeepSeek-R1-Distill series. After examination, all annotators confirmed that the outputs of LRMs contain no harmful content, resulting in $\mathbf{100\%}$ inter-annotator agreement. We provide additional examples of Whisper in Figures 10 to 14. These examples illustrate that, compared to the standard CoT prompt, Whisper substantially mitigates overthinking in LRMs and elicits concise reasoning without compromising performance.
Figure 10: An illustrated case of Whisper on DeepSeek-R1-Distill-LLaMA-8B under evidence-based persuasion.
Figure 11: An illustrated case of Whisper on DeepSeek-R1-Distill-Qwen-14B under role-playing persuasion.
Figure 12: An illustrated case of Whisper on Qwen3-14B, under the persuasion perspective of emotional appeal.
Figure 13: An illustrated case of Whisper on Qwen3-14B under threat-based persuasion.
Figure 14: An illustrated case of Whisper on commonsense reasoning. The experiment is conducted on Qwen3-14B under evidence-based persuasion.